# World In Your Hands: A Large-Scale and Open-source Ecosystem for Learning Human-centric Manipulation in the Wild

### v1

@@ Abstract
Large-scale pre-training is fundamental for generalization in language and vision models, but data for dexterous hand manipulation remains limited in scale and diversity, hindering policy generalization. Limited scenario diversity, misaligned modalities, and insufficient benchmarking constrain current human manipulation datasets. To address these gaps, we introduce World In Your Hands (WiYH), a large-scale open-source ecosystem for human-centric manipulation learning. WiYH includes (1) the Oracle Suite, a wearable data collection kit with an auto-labeling pipeline for accurate motion capture; (2) the WiYH Dataset, featuring over 1,000 hours of multi-modal manipulation data across hundreds of skills in diverse real-world scenarios; and (3) extensive annotations and benchmarks supporting tasks from perception to action. Furthermore, experiments based on the WiYH ecosystem show that integrating WiYH’s human-centric data significantly enhances the generalization and robustness of dexterous hand policies in tabletop manipulation tasks. We believe that World In Your Hands will bring new insights into human-centric data collection and policy learning to the community.
@@ 1 Introduction
| Dataset | Hours | Clips | Wild | Skills | Cam. Calib | RGB | Tactile | Action | Depth | Mask | Instruct. | Reason |
| Ego4D [grauman2022ego4d] | 3,670 | - | ✓ | ✓ | ✗ | ✓ | ✗ | 2D | ✗ | ✗ | ✓ | ✓ |
| HOI4D [liu2022hoi4d] | 44.4 | 4k | ✓ | ✓ | ✓ | ✓ | ✗ | 2D | ✓ | ✓ | ✗ | ✗ |
| HoloAssist [wang2023holoassist] | 166 | 2.22k | ✗ | ✗ | ✗ | ✓ | ✗ | - | ✓ | ✗ | ✓ | ✓ |
| Ego-Exo4D [grauman2024ego] | 1,286 | 5.04k | ✓ | ✓ | ✓ | ✓ | ✗ | ✗ | ✗ | ✓ | ✓ | ✓ |
| CaptainCook4D [peddi2024captaincook4d] | 94.5 | 384 | ✓ | ✓ | ✗ | ✓ | ✗ | ✗ | ✓ | ✗ | ✓ | ✓ |
| HOT3D [banerjee2025hot3d] | 13.9 | 425 | ✗ | ✗ | ✓ | ✓ | ✗ | 3D | ✗ | ✗ | ✗ | ✗ |
| HO-Cap [wang2024ho] | - | 64 | ✗ | ✗ | ✓ | ✓ | ✗ | 3D | ✓ | ✓ | ✗ | ✗ |
| Ego-ViD-5M [wang2024egovid] | - | 5M | ✓ | ✗ | ✓ | ✓ | ✗ | ✗ | ✗ | ✗ | ✓ | ✗ |
| EgoDex [hoque2025egodex] | 829 | 338k | ✗ | ✓ | ✗ | ✓ | ✗ | 3D | ✗ | ✓ | ✓ | ✗ |
| WiYH (Ours) | 1045 | 125.4k | ✓ | ✓ | ✓ | ✓ | ✓ | 3D | ✓ | ✓ | ✓ | ✓ |
Large-scale pre-training has been established as a cornerstone for achieving generalization in Large Language Models [brown2020language, chowdhery2023palm, achiam2023gpt, touvron2023llama, yang2025qwen3], Vision-Language Models [radford2021learning, li2022blip, alayrac2022flamingo, li2023blip, li2024llava] and Vision-Language-Action Models [brohan2022rt, zitkovich2023rt, kim2024openvla, black2410pi0, wen2024diffusion, zhou2025vision]. However, datasets for dexterous hand manipulation remain far smaller and less diverse than language datasets, limiting progress in learning robust manipulation policies.
To address the problem of data scarcity, human-centric data collection and learning have attracted growing interest [hsieh2025dexman, zhu2025learning, routray2025vipra, zhu2025let, kareer2025egomimic]. Some approaches [chao2021dexycb, kwon2021h2o, banerjee2025hot3d] capture human hand data during object manipulation using devices such as VR systems and smart glasses, while others [grauman2022ego4d, xu2021videoclip, wang2021actionclip, he2024egovm] leverage foundation models to extract human actions from unlabeled videos of human manipulation available on the internet.
Despite these advances, current human manipulation datasets still face three major limitations, as shown in Table 1 that hinder their usefulness for dexterous manipulation research: (1) Limited Scenario Diversity: Many datasets are collected in constrained laboratory environments, lacking the diversity and complexity of real-world settings. (2) Inadequate Alignment: Existing collections often lack one or more key modalities, such as the absence of fine-grained language instructions, precise 3D hand poses, or object poses, resulting in misalignment across vision, language, and action. (3) Insufficient Benchmarking: Most datasets do not support comprehensive evaluation benchmarks that span the entire pipeline of dexterous manipulation, including scene perception, task planning, and action.
@@
The Oracle Suite and Auto-labeling Pipeline. We introduce a lightweight and flexible wearable data collection kit, the Oracle Suite, designed to capture data without impeding the operator’s natural movements. An integrated auto-labeling pipeline enables markerless self-localization and centimeter-level motion accuracy without requiring external third-person tracking.
The WiYH Dataset: A Large-Scale Multi-modal Human-centric Dataset. We present the WiYH dataset, a large-scale multi-modal resource comprising over 1,000 hours of manipulation sequences across hundreds of skills, captured in 10 diverse industrial and daily-life scenarios such as factories, supermarkets, hotels, and restaurants. The dataset includes rich sensory streams: multi-view images, camera calibration parameters, depth maps, 3D hand poses, and 6D wrist trajectories.
Diverse Annotations and Comprehensive Benchmarks. Building upon the WiYH dataset, we provide extensive annotations, including approximately 400 hours of atomic-level action instructions and 100k instances of vision-language data for task reasoning, affordance understanding, and object localization. These annotations support a comprehensive set of benchmarks—spanning visual language understanding, world model evaluation, and vision-language-action tasks—designed to assess capabilities from scene perception to embodied operation holistically.
Based on the WiYH ecosystem, we study how human-centric data empowers dexterous manipulation learning in tabletop manipulation tasks. We train models on mixtures of Oracle Suite demonstrations and other manipulation data and evaluate the task success rate across varied environments and object layouts. Our experiments demonstrate that human-centric data substantially improves the generalization and robustness of dexterous hand policies.
[Limitation]. All manipulation tasks are evaluated in real-world environments. To increase accessibility and reproducibility, we are developing high-fidelity simulation environments and plan to release them in a future version of the dataset and benchmarks.
@@ 2.1 Oracle Suite: Human-centric Data Collection Suite
To enable large-scale, high-fidelity, human-centric multimodal data acquisition, we developed Oracle Suite, a low-cost wearable data collection system designed for human-centric embodied AI research. Specifically, as shown in Figure 2, Oracle Suite comprises a hardware system for acquiring multi-view RGB, IMU-based localization, and tactile information, alongside an autolabeling module that achieves high-precision motion capture through the fusion of IR, RGB, and IMU localization.
@@
Hardware System Overview. Oracle Suite adopts a modular architecture integrating three primary hardware components: H-HPVHive, H-Gloves, and H-Backpack. This modular design enables flexible deployment across diverse environments, without being restricted to controlled indoor labs.
@@
The Oracle Suite introduces major optimizations in ergonomics, sensing, and reliability:
@@
H-FPVHive: It supports chest-mounted configuration, and integrates two fisheye cameras, two pinhole cameras, and four infrared lenses. All cameras are hardware synchronized. The infrared lenses are used for better locating the relative poses of H-Gloves.
@@
H-Gloves: Each glove contains 6 IMUs, 5 pressure sensors, and an on-board MCU. The gloves can capture the hand motions and tactile. Moreover, it also contains 3 fisheye cameras for visual tracking and observation.
@@
H-Backpack: It provides data storage, computation, and power supply. The computation is hosted by NVIDIA Orin.
Autolabeling Pipeline. As shown in the right part of Figure 2, the Autolabeling Pipeline comprises both online and offline algorithms. It fuses multi-modal sensor data (IR + IMU + RGB) to ensure positioning accuracy in in-the-wild scenarios. Specifically, the online localization module performs lightweight coarse positioning using visual-inertial odometry (VIO) algorithms and IR detection, while the offline stage refines these coarse estimates through visual Structure-from-Motion (SfM) algorithms. The pipeline ultimately outputs centimeter-level 6D poses for both wrists.
Data Production Efficiency. The Oracle Suite pipeline follows a collection–annotation–inspection–archiving workflow for scalable data production. A single collector operating 5 hours per day generates approximately 1.8 TB of valid multimodal data. At scale, 100 collectors over 30 days yield roughly 5.4 PB of raw data (stored in cloud repositories). Efficiency is further improved through standardized capture protocols and automated cloud-based annotation.
2.2 Data Annotation Pipeline
@@
Data collection: Practitioners wore the Oracle Suite in each scene to record natural manipulation behavior. Each recording session lasted approximately 10 minutes and included start and end event markers. The collected signals comprise tactile signals from pressure sensors, binocular fisheye images from the left and right wrist mounts, first-person binocular fisheye images, left-hand and right-hand IMU localization data, and the corresponding camera calibration parameters. Samples of data annotations are shown in Figure 4.
Data annotation: The annotation pipeline comprises three parts: subtask (atomic task) annotation, perception annotation, and vision-language annotation.
Subtask annotation: To construct a large-scale, high-quality set of atomic tasks, we designed a four-stage hybrid annotation pipeline consisting of data processing, task definition, annotation consolidation, and post-processing:
(1) Data processing: First-person fisheye videos are segmented into 30-second clips, which serve as the basic units for semantic annotation by the video-language module;
(2) Task definition (manual annotation): The video-language module parses hierarchical operational semantics from the first-person videos and hierarchically decomposes an operation sequence into high-level tasks and atomic subtask units. Each task unit contains a preset action described in natural language and a manually annotated target object.
(3) Data checking (manual review): Outputs from the video-language module are subject to iterative manual review; any incorrect annotations are reprocessed before proceeding to the consolidation stage.
(4) Post-processing: Annotated preset-action and target-object names are assembled into complete operation instructions and are further augmented using a large language model (LLM). After subtask annotation, the raw data are organized into a hierarchical format of tasks composed of multiple atomic subtasks.
Perception annotation: To provide richer perceptual information, we pre-annotated scene elements using visual foundation models and then performed manual selection and correction. The resulting perception annotations include:
(1) Segmentation masks: instance masks are generated by applying SAM to human-provided prompt points;
(2) Depth maps: scene depth is estimated from first-person binocular images using the FoundationStereo model.
Vision-language annotation: The vision-language annotations were produced through manual curation and comprise the following tasks:
(1) Spatial Referring: For each subtask, we extract a key frame near the action start that contains the operator’s object of attention and in which the target object is not occluded by the hand, body, or other objects. Based on the subtask’s target object, spatial relational terms are added to uniquely refer to a single target region in the image.
(2) Subtask Prediction: We select tasks that contain 4–5 subtasks. The next subtask following the current one is used as the correct option. Nine incorrect options are generated by combining scene object information and the task’s preset actions; annotators then select the three incorrect options that are the most logically plausible. Additionally, Qwen-2.5VL-72B was used to generate auxiliary “thinking” data for next-action prediction, which were subsequently corrected by human annotators;
(3) Completion Verification: We select tasks containing 4-5 subtasks and randomly trim the tail of the video by 20%, 25%, 30%, or 35% of frames. The task description and the truncated video are provided as input, and human annotators label whether the video depicts a completed task.
Through the above procedures, we obtain rich annotations covering scene understanding and task planning, which constitute the foundation of a unified benchmark.
@@ 2.3 Data Statistics
Scenes such as supermarkets and laundry include long-horizon activities with substantial execution time, while hotels and candlelight contain shorter but highly structured routines. This diversity in task–scene composition highlights the wide operational range covered by WiYH. Action-level temporal characteristics are shown in Figure 3(b). For each scene, we compute the average execution duration of fine-grained actions. The dataset exhibits substantial variation across action categories, from short, instantaneous operations such as wipe and insert, to long-horizon manipulation such as smooth out or foldlay. These temporal differences reflect the heterogeneous complexity of real-world household and industrial workflows.
@@
To better understand the manipulation space in WiYH, Figure 3(d–e) visualizes word clouds of the annotated target objects and skills. The dataset covers a broad range of manipulable objects—including clothes, linens, cartons, cups, and tableware—spanning deformable, rigid, and articulated categories. Likewise, the skill vocabulary encompasses diverse manipulation primitives such as take, place, rotate, unfold, align, push, and search, representing both low-level motor actions and higher-level task strategies. These distributions further demonstrate that WiYH captures a wide spectrum of real-world manipulation behaviors.
3 Benchmark and Applications
3.1 Human-centric Vision-Language (HVL)

### v3

@@ Abstract
We introduce World In Your Hands (WIYH), a large-scale open-source ecosystem comprising over 1,000 hours of human manipulation data collected in-the-wild with millimeter-scale motion accuracy. Specifically, WIYH includes (1) the Oracle Suite, a wearable data collection kit with an auto-labeling pipeline for accurate motion capture; (2) the WIYH Dataset, featuring over 1,000 hours of multimodal manipulation data across hundreds of skills in diverse real-world scenarios; and (3) extensive annotations and benchmarks supporting tasks from perception to action. Furthermore, experiments based on the WIYH ecosystem show that integrating WIYH’s human-centric data improves robotic manipulation success rates from 8% to 60% in cluttered scenes. World In Your Hands provides a foundation for advancing human-centric data collection and cross-embodiment policy learning. All data and hardware design will be open-source.
@@ 1 Introduction
Large-scale pre-training has been established as a cornerstone for achieving generalization in Large Language Models [brown2020language, chowdhery2023palm, achiam2023gpt, touvron2023llama, yang2025qwen3], Vision-Language Models [radford2021learning, li2022blip, alayrac2022flamingo, li2023blip, li2024llava] and Vision-Language-Action (VLA) Models [brohan2022rt, zitkovich2023rt, kim2024openvla, black2410pi0, wen2024diffusion, zhou2025vision]. However, datasets for dexterous hand manipulation remain far smaller and less diverse than language datasets, limiting progress in learning robust manipulation. To address the problem of data scarcity, human-centric data collection and learning have attracted growing interest [hsieh2025dexman, zhu2025learning, routray2025vipra, zhu2025let, kareer2025egomimic]. Some approaches [chao2021dexycb, kwon2021h2o, banerjee2025hot3d] capture human hand data during object manipulation using devices such as VR systems and smart glasses, while others [grauman2022ego4d, xu2021videoclip, wang2021actionclip, he2024egovm] leverage foundation models to extract human actions from unlabeled videos of human manipulation available on the internet.
| Dataset | Hours | Clips | Wild | Cam. Calib | RGB | Tactile | Action | Depth | Mask | Instruct. | VLM |
| Ego4D [grauman2022ego4d] | 3,670 | ✓ | ✓ | ✗ | ✓ | ✗ | 2D | ✗ | ✗ | ✓ | ✓ |
| HOI4D [liu2022hoi4d] | 44.4 | 4k | ✓ | ✓ | ✓ | ✗ | 2D | ✓ | ✓ | ✗ | ✗ |
| HoloAssist [wang2023holoassist] | 166 | 2.22k | ✗ | ✗ | ✓ | ✗ | - | ✓ | ✗ | ✓ | ✓ |
| Ego-Exo4D [grauman2024ego] | 1,286 | 5.04k | ✓ | ✓ | ✓ | ✗ | ✗ | ✗ | ✓ | ✓ | ✓ |
| CaptainCook4D [peddi2024captaincook4d] | 94.5 | 384 | ✓ | ✗ | ✓ | ✗ | ✗ | ✓ | ✗ | ✓ | ✓ |
| HOT3D [banerjee2025hot3d] | 13.9 | 425 | ✗ | ✓ | ✓ | ✗ | 3D | ✗ | ✗ | ✗ | ✗ |
| HO-Cap [wang2024ho] | - | 64 | ✗ | ✓ | ✓ | ✗ | 3D | ✓ | ✓ | ✗ | ✗ |
| Ego-ViD-5M [wang2024egovid] | - | 5M | ✗ | ✓ | ✓ | ✗ | ✗ | ✗ | ✗ | ✓ | ✗ |
| EgoDex [hoque2025egodex] | 829 | 338k | ✓ | ✗ | ✓ | ✗ | 3D | ✗ | ✓ | ✓ | ✗ |
| WIYH (Ours) | 1045 | 125.4k | ✓ | ✓ | ✓ | ✓ | 3D | ✓ | ✓ | ✓ | ✓ |
Despite these advances, current human manipulation datasets still face three major limitations, as shown in Table 1 that hinder their usefulness for dexterous manipulation research: (1) Limited Scenario Diversity: Many datasets are collected in constrained laboratory environments, lacking the diversity and complexity of real-world settings. (2) Inadequate Alignment: Existing collections often lack one or more key modalities, such as the absence of fine-grained language instructions, precise 3D action, or object poses, resulting in misalignment across vision, language, and action. (3) Insufficient Benchmarking: Most datasets do not support comprehensive evaluation benchmarks that span the entire pipeline of dexterous manipulation, including scene perception, task planning, and action.
@@
The Oracle Suite. We introduce a lightweight, flexible wearable data-collection kit that captures 3D action data with millimeter-scale translational accuracy (mean error below 5 mm) without external third-person tracking.
The WIYH Dataset. We present a large-scale multi-modal resource comprising over 1,000 hours of manipulation demonstrations across 100 skills, captured in 10 diverse scenarios. The dataset includes rich sensory streams: multi-view images, camera calibration parameters, hand poses, and wrist trajectories.
Diverse Annotations. Building upon the WIYH dataset, we provide extensive annotations, including 400 hours of atomic action instructions and 100k instances of vision-language data, supporting a comprehensive usage from scene perception to robotic manipulation.
Based on the WIYH ecosystem, we study how human-centric data empowers dexterous manipulation learning in robotic manipulation tasks. We conduct two sets of experiments: (1) cross-embodiment pretraining and (2) retargeting co-training, and evaluate the task success rate across varied environments and object layouts. Our experiments demonstrate that human-centric data substantially improves the generalization and robustness of robotic manipulation, whether as large-scale cross-embodiment pre-training data or retargeted action data.
Limitation. To increase accessibility and reproducibility, we are developing high-fidelity simulation environments and plan to release them in a future version of the dataset and benchmarks.
@@ 2.1 Oracle Suite: Human-centric Data Collection Suite
We develop Oracle Suite, a low-cost wearable data collection system designed for human-centric multimodal data collection. Specifically, as illustrated in Figure 2, Oracle Suite comprises a hardware system that captures multi-view RGB images, IMU-based localization, and tactile information and an autolabeling module that achieves high-precision motion capture by fusing IR, RGB, and IMU localization.
@@
Hardware System Overview. Oracle Suite adopts a modular architecture integrating three primary hardware components: H-FPVHive, H-Gloves, and H-Backpack. This design enables flexible deployment across diverse environments without being restricted to stable indoor labs.
@@
Hardware Design. The Oracle Suite incorporates significant optimizations in ergonomics, sensing, and reliability.
@@
H-FPVHive supports a chest-mounted configuration with two fisheye cameras, two pinhole cameras, and four infrared lenses integrated into the chest unit, where all cameras achieve hardware synchronization, and the infrared lenses facilitate better localization of the relative pose of H-Gloves.
@@
H-Gloves include six IMUs, five fingertip pressure sensors with a resolution of 5mN and a range of 0.2 N to 50 N, and an onboard MCU per glove to capture hand motions and tactile information while also containing three fisheye cameras for visual tracking and observation.
@@
H-Backpack provides data storage, computation, and power supply, utilizing an NVIDIA Orin computing core.
Autolabeling Pipeline. The autolabeling pipeline involves online and offline algorithms, as shown on the right side of Figure 2. By fusing multimodal sensor streams (IR, IMU, and RGB), it maintains robust localization in in-the-wild scenarios and ultimately outputs 6D wrist pose trajectories with millimeter-scale translational accuracy (mean positional error below 5 mm).
Efficiency of Oracle Suite. Compared to teleoperation, Oracle Suite achieves 5 higher collection efficiency (720 vs. 150 episodes/day) without extensive operator training. Compared to VR-based hand tracking, Oracle Suite provides more accurate 3D hand skeletons, particularly under visual occlusion.
2.2 Data Validation System
We employ motion capture tests and projection intersection methods to validate data quality. Figure 3 illustrates the testing effects of both methods, and we retain valid high-quality data. For the motion capture test, we attach motion capture spheres to H-Gloves and complete operations identical to those in the WIYH dataset within a motion-capture room. This setup tests the deviation between the wrist trajectory generated by the Oracle Suite autolabeling pipeline and the motion capture trajectory to verify the effectiveness of the automatic annotation pipeline. Our Oracle Suite achieves an average position error of 5 mm in the motion capture room environment. For projection overlap, we obtain hand masks in perspective views of H-FPVHive using a hand segmentation algorithm. We then project the skeleton data from H-Gloves into the same view and calculate the intersection with the hand mask to determine accuracy.
3 World In Your Hands Dataset
3.1 Data Collection and Annotation
@@
Data collection: During data collection, collectors wear the Oracle Suite in various scenarios to record natural manipulation behaviors. Each task is recorded once according to the Standard Operating Procedure of the scenario. Recording sessions last approximately 10 minutes and include start and end event markers. Captured signals include tactile signals from pressure sensors, wrist images, first-person chest images, corresponding camera calibration parameters, left and right hand trajectory data, and left and right hand skeleton information. Samples of data annotations are shown in Figure 6.
After data collection, we curate the WIYH dataset via a three-stage annotation pipeline (Figure 4), comprising atomic action annotation, perception annotation, and vision-language annotation.
Atomic action annotation: To construct data suitable for robotic foundation model training, we design a four-stage hybrid annotation process for atomic action annotation that decomposes a task in the standard operating procedure into multiple atomic segments, where each segment is termed a subtask.
Video Segmentation: The process initiates with video segmentation, where Qwen2.5-VL-72B [Qwen2.5-VL] segments the recorded data for each task into semantic atomic fragments serving as basic units for annotation.
Task definition: Humans define a preset action termed a "skill" and an operation target object for each atomic fragment.
Data checking: Manual reviewers check boundary and task definition accuracy to correct errors.
Instruction Augmentation: Post-processing combines the annotated preset actions and target object names into complete atomic instructions, which are augmented and enhanced using Qwen2.5-VL-72B [Qwen2.5-VL].
Perception annotation: We use vision foundation models [kirillov2023segany, wen2025foundationstereo] to pre-annotate scene elements, followed by manual selection and correction. This includes generating instance masks by applying SAM [kirillov2023segany] to human-provided prompt points and estimating scene depth from first-person binocular images using the FoundationStereo [wen2025foundationstereo] model.
Vision-language annotation is generated through manual curation and includes several components.
CoT Reasoning: We use Qwen2.5-VL-72B [Qwen2.5-VL] to generate a thought process indicating the position of each subtask within a complete task and when to enter the next subtask for long-horizon task decomposition, where generated thought data is reviewed and corrected by humans.
Spatial Referring (SR): We select key frames near the action start where the operation target object is clearly visible for some subtasks and add spatial relation terms based on the atomic instruction to uniquely refer to a single target area in the scene.
Subtask Prediction (SP): We select tasks containing 4 to 5 subtasks and use the next subtask after the current one as the correct option, while generating three logically reasonable incorrect options based on the task description.
Completion Verification (CV): We randomly crop 0% to 35% of each subtask for a task and use the atomic instruction of each subtask, as well as the cropped fragment, as input to label the task progress status regarding whether the task is completed.
Notably, subtask prediction, spatial referring, and progress verification are annotated on only a subset of the whole WIYH data to construct an evaluation set for validating the capabilities of existing VLMs in human-centric scenarios.
@@ 3.2 Data Statistics
Scenes such as supermarkets and laundry include long-horizon activities with substantial execution time, while hotels and candlelight contain shorter but highly structured routines. This diversity in task–scene composition highlights the wide operational range covered by WIYH. Action-level temporal characteristics are shown in Figure 5(b). For each scene, we compute the average execution duration of fine-grained actions. The dataset exhibits substantial variation across action categories, from short, instantaneous operations such as "wipe" and "insert" to long-horizon manipulation such as "smooth out" or "fold lay". These temporal differences reflect the heterogeneous complexity of real-world household and industrial workflows.
@@
To better understand the manipulation space in WIYH, Figure 5(d–e) visualizes word clouds of the annotated target objects and skills. The dataset covers a broad range of manipulable objects—including clothes, linens, cartons, cups, and tableware—spanning deformable, rigid, and articulated categories. Likewise, the skill vocabulary encompasses diverse manipulation primitives such as take, place, rotate, unfold, align, push, and search, representing both low-level motor actions and higher-level task strategies. These distributions further demonstrate that WIYH captures a wide spectrum of real-world manipulation behaviors. More data, demonstrations, and statistics will be provided in the supplementary materials.
4 Benchmark and Application
4.1 Benchmark: Human-centric Vision-Language (HVL)